# Symmetry-broken charge-ordered ground state in CsV3Sb5 Kagome metal

### v2

@@ Abstract
By calculating the phonon spectral function with the dynamical extension of the SSCHA theory [46, 48, 49], we predict that the softening of the $T_{1}^{+}$ mode should be observable by, for instance, inelastic X-ray experiments. As shown in Fig. 4(e), this mode undergoes a significant renormalization before collapsing at 80 K, but it remains separated from the rest of the spectra, allowing its spectral peak to remain well-defined still at 75 K. We perform a similar analysis for the $\Gamma_{1}^{+}$ phonon mode, the one that emerges as the condensation of the $L_{2}^{-}$ mode of the high-symmetry phase. This mode exhibits a marked red-shift from 93 to 78 cm-1 between 0 and 75 K, accompanied by an increase in broadening, which makes it practically vanish from the spectrum at 75 K (see Fig. 4(d)). This behavior aligns well with previous Raman measurements, which report a strong renormalization of a $\Gamma_{1}^{+}$ mode at similar energies [52].
@@
In summary, by performing first-principles calculations including the ionic kinetic energy, entropy, and anharmonicity, we resolve the phase diagram of CsV3Sb5 and its CDW phase. Our results yield a first-order CDW phase transition at 89 K, in good agreement with experimental results. The obtained CDW ground-state ordering is inherently three dimensional, with the nature of the stacking being hard to resolve. However, being consistent with transport and Raman data, the obtained $\pi$-TrH and 4TrH phases are realistic models of the CDW, and offer new avenues to the understanding of its still mysterious superconductivity and time-reversal symmetry breaking.
@@ Acknowledgements.
This project was supported by the Spanish Ministerio de Ciencia e Innovación (grants PID2022-142008NB-I00 and PID2022-142861NA-I00). I.E. acknowledges the Department of Education, Universities and Research of the Eusko Jaurlaritza and the University of the Basque Country UPV/EHU (Grant No. IT1527-22). M.G.-A. was supported by Jane and Aatos Erkko Foundation, Keele Foundation, and Magnus Ehrnrooth Foundation as part of the SuperC collaboration. M.A. acknowledges a PhD scholarship from the Materials Physics Center. The author acknowledges the technical and human support provided by the DIPC Supercomputing Center.

### v3

@@ Abstract
By calculating the phonon spectral function with the dynamical extension of the SSCHA theory [50, 53, 54] along with first-principle electron-phonon calculations, we predict that the softening of the $T_{1}^{+}$ mode should be observable by, for instance, inelastic X-ray experiments. As shown in Fig. 4(d), this mode undergoes a significant renormalization before collapsing at 80 K, but it remains separated from the rest of the spectra, allowing its spectral peak to remain well-defined still at 75 K. We perform a similar analysis for the $\Gamma_{1}^{+}$ phonon mode, the one that emerges as the condensation of the $L_{2}^{-}$ mode of the high-symmetry phase. This mode exhibits a marked red-shift from 93 to 78 cm-1 between 0 and 75 K, accompanied by an increase in broadening, which makes it practically vanish from the spectrum at 75 K. This behavior aligns well with previous Raman measurements, which report a strong renormalization of a $\Gamma_{1}^{+}$ mode at similar energies [58].
Having determined the charge-ordered structure allows us to study the origin of the intriguing superconductivity of CsV3Sb5 from first principles. Experimentally, the superconducting critical temperature ($T_{c}$) of approximately 2.5 K [14] increases under pressure displaying a double-dome structure before superconductivity is ultimately suppressed [59, 60]. The first of these two $T_{c}$ maxima coincides with the previously discussed emergence of the in-plane superlattice phase, while the second corresponds to the melting of the CDW, where $T_{c}$ increases approximately up to 8 K [56]. Despite the electron-phonon interaction was estimated to be too weak to explain the observed $T_{c}$ [18], pointing to an unconventional mechanism [17, 18, 19], the phonon mediated scenario cannot be ruled out since no calculation has been performed thus far in the CDW phase considering the crucial role of anharmonicity. By combining the SSCHA anharmonic phonon spectrum of the $\pi$-TrH phase with ab initio electron-phonon matrix elements (see Supplementary Material), we calculate a $T_{c}$ of 2.7 K for the CDW phase, in good agreement with experiments. As shown in Fig. 4(e), the contribution to the electron-phonon coupling constant $\lambda$ is rather evenly distributed among all phonon modes. In order to test the impact of the CDW on the superconducting critical temperature, we repeat the calculation for the high-symmetry phase assuming that its phonon spectra is the anharmonic one at 100 K. Interestingly, $T_{c}$ is enhanced to 12.1 K, which is consistent with the large increase of the critical temperature observed experimentally when the CDW is melted under pressure. In fact, the CDW strongly suppresses the electron-phonon coupling constant (see Suppementary Material). Our results demonstrate that superconductivity in CsV3Sb5 is conventional, in the sense that it is mediated by the electron-phonon coupling, and that there is a large interaction between superconductivity and the CDW reconstruction.
@@
In summary, by performing first-principles calculations including the ionic kinetic energy, entropy, and anharmonicity, we resolve the phase diagram of CsV3Sb5 and its CDW phase. Our results yield a first-order CDW phase transition at 89 K, in good agreement with experimental results. The obtained CDW ground-state ordering is inherently three dimensional, with the nature of the stacking being hard to resolve. However, being consistent with transport and Raman data, the obtained $\pi$-TrH and 4TrH phases are realistic models of the CDW. With the CDW solved and having a realistic description of its phonon spectrum, we determine that superconductivity is mediated by phonons and that the CDW strongly suppresses its critical temperature.
@@ Acknowledgements.
We thank Fernando de Juan for insightful discussions. This project was supported by the PID2022-142861NA-I00 and PID2022-142008NB-I00 projects funded by MICIU/AEI/10.13039/501100011033 and FEDER, UE; the Simons Foundation through the Collaboration on New Frontiers in Superconductivity (Grant No. SFI-MPS-NFS-00006741-10); the Department of Education, Universities and Research of the Eusko Jaurlaritza and the University of the Basque Country UPV/EHU (Grant No. IT1527-22); Canada Excellence Research Chairs Program for Topological Quantum Matter; NSERC Quantum Alliance France-Canada; and Diputación Foral de Gipuzkoa Programa Mujeres y Ciencia M.A. acknowledges a PhD scholarship from the Materials Physics Center. M.G.-A. was supported by Jane and Aatos Erkko Foundation, Keele Foundation, and Magnus Ehrnrooth Foundation as part of the SuperC collaboration. The authors acknowledge the technical and human support provided by the DIPC Supercomputing Center.